OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-2, 7) and (-7, 12). Please show all of your work.

OpenStudy (anonymous):

first find slope \[\Large m=\frac{12-7}{-7-(-2)}=\frac{5}{-5}=-1\] now use point slope form \[\Large y-y_{0}=m(x-x_{0})\] where you can use any point (-2, 7) OR (-7, 12) then express your answer in the form y=mx+c

OpenStudy (anonymous):

@sharron482 can you do this now ?

OpenStudy (anonymous):

i think im good now, thank you